# Distributed network of smartphone sensors: a new tool for scientific field measurements

### v2

@@ Abstract
Smartphones sensors are now commonly used by a worldwide audience thanks to their availability, high connectivity, and versatility. Here, we present a methodology to use a collection of smartphones, namely a fleet, as a distributed network of time-synchronized mechanical sensors. We first present the mechanical tests we develop to evaluate the smartphone sensor accuracy. We then describe how to use efficiently a distributed network of smartphones as autonomous sensors. We use a combination of an Android application hosted on each phone (Gobannos), and a master application (Phonefleet) from a controlling host to perform the tasks in parallel remotely. We detail the time synchronisation protocol we implemented, based on UDP communication. We achieved an accuracy of the time synchronisation of 50 micro-seconds, better enough for all mechanical applications. Eventually, we present two test cases in which we use a smartphone fleet for mechanical wave measurements in field conditions.
preprint: AIP/123-QED
I Introduction
@@
By November 2024, mobile subscriptions have reached 8.4 billion, including 7.14 billion for smartphones Jonsson (2024). Smartphones have become a key tool for educational and scientific research purposes thanks to their accessibility to the general public Gao and Wu (2016); Raju et al. (2024) and to the versatility of their built-in micro-electro-mechanical system (MEMS) sensors. A typical MEMS is a highly integrated, silicon-based sensor system fabricated using semiconductor processes Maluf (2002). Its low cost, compact size, lightweight design, and low power consumption make a MEMS ideal for integration into smartphones, drones, tablets, and wearable devices Daponte et al. (2013), enabling applications such as step counting and orientation tracking Bao and Intille (2004); Naqvib et al. (2012); Davidson and Piché (2016), gaming Höpfner et al. (2013); Kołakowska, Szwoch, and Szwoch (2020), or navigation Li et al. (2015); Mostafa et al. (2019).
@@
In the context of physics education, following the COVID-19 pandemic, smartphones have gained popularity as flexible tools for providing innovative teaching methodologies, particularly at the university level Bobroff, Bouquet, and Delabre (2020); Bouquet et al. (2020); O’Brien (2021). Numerous examples of tabletop physics experiments using smartphone as a measurement device can be found, such as speed of sound measurement using the embedded microphone Kasper, Vogt, and Strohmeyer (2015); earth’s rotation estimate using the accelerometer Vandermarlière (2021); and light polarization analyses using ambient light sensor Monteiro et al. (2017). A comprehensive reviews of applications can be found in Organtini Organtini et al. (2021) or in Kuhn &\&& Vogt  Kuhn and Vogt (2022).
@@
The versatility of smartphone MEMS sensors has also drawn significant attention from researchers across various scientific domains. The high connectivity and ubiquity of smartphones enable large-scale, real-time data collection through a geographically distributed network of interconnected devices. For public health research, investigations on the data collection and analysis of energy expenditure from smartphone users Pande et al. (2013) have shown that smartphone-embedded IMU (Inertial Measurement Unit) sensors such as accelerometers and gyroscopes, can accurately monitor the user’s physical activities, thus enhance the precision of individual health assessments. In a review by Lee et al. Lee, Suh, and Choi (2018) on smartphone applications in geoscience research, the authors highlight the growing use of smartphones in tasks traditionally carried out with specialized tools, such as geological mapping, seismic activity detection, and natural hazard assessment.
@@
However, field measurements often require multi-point data collection with sensors spatially distributed over a wide area while remaining synchronized in time. Examples of such applications include predicting solar magnetic storms and assessing their impact on Earth’s magnetosphere Odenwald (2022), as well as performing modal analysis of civil engineering structures Guéguen et al. (2020), where a large number of data points are crucial for reconstructing higher-order spatial modes Cunha and Caetano (2006); Rainieri and Fabbrocino (2014). Distributed sensor measurements in seismology using smartphones have been introduced via real-time data collections for earthquake detection using the application MyShake Kong, Allen, and Schreier (2016). It allows accurate localization of the earthquake’s epicenter and detects seismic waves propagation in real time, contributing to early warning systems Reilly et al. (2013); Kong, Allen, and Schreier (2016). These applications do not rely on sophisticated, high-cost sensors but rather utilize frugal, cost-effective sensor networks, such as smartphone-based MEMS sensors fleets on a large scale, often in regions geographically inaccessible through traditional methods.
@@
Several mobile applications (PhyPhox, FizziQStaacks et al. (2018); Bilgin, Molina Ascanio, and Minoli (2022) have been developed for individual users to facilitate sensor communication and to display data in real-time. However, there is a notable lack of a technical way to control multiple smartphones by a single user. In contrast to single smartphone measurement that can be performed with manual interactions, distributed measurements require an automated process involving remote communication for data sampling and collection, time synchronization between smartphones, and parallel task planification. Another key challenge lies in the optimization of human-machine interaction (HMI) for a single user to control a fleet smartphone.
In this study, we present a methodology to perform simultaneous multi-sensor measurements using a fleet of identical smartphones. The article is structured as follows: we first review the measurement accuracy of a single smartphone using several test set up. We characterize in particular the ambient noise amplitude and spectrum of a single smartphone placed in a still environment. We then perform accelerometer and gyroscope sensor calibration on a turntable platform Mau et al. (2016) and we determine the sensor location within the phone. This retro-engineering approach allows to access information that are otherwise proprietary for most commercial products. Next, we develop an automated remote communication and control protocol for data recording across the smartphone fleet, achieving time synchronization with a typical error of 50 $\mu$s. To do so, we introduce a custom Android-based application, Gobannos Casoli et al. (2025), which integrates memory allocation for local data storage, remote communication, and time synchronization. We eventually provide two examples of mechanical wave measurements, demonstrating the feasibility of reliable data collection using a large fleet of smartphones.
II Smartphones as a multi-sensor fleet
@@
To create a fleet of autonomous, multi-physics sensors, we selected the Redmi 10A smartphone, balancing the cost and the quality of the embedded IMU sensors (accelerometer, gyroscope and magnetometer). However, the methodology descrived in this manuscript may apply to all types of smartphones. The fleet is composed of 66 smartphones, numbered from 0 to 65. We use concomitantly the accelerometer, gyroscope, magnetometer and GPS sensors. To access and utilize the smartphone sensors, we initially employed the Phyphox application Staacks et al. (2018), developed at the University of Aachen. Phyphox offers an intuitive interface for acquiring data from individual sensors or from multiple sensors simultaneously, along with a standard URL communication protocol to retrieve data from the smartphone. We configured a module in Phyphox to simultaneously collect accelerometer, gyroscope, magnetometer, and GPS data at their respective maximum sampling frequencies. In a second step, we develop an Android application, named Gobannos Casoli et al. (2025), better adapted to parallelized, continuous and time synchronized acquisitions without physical intervention on the phones. This section is organized as follow. Section A presents the mechanical tests performed on the smartphone sensors. Section B details the architecture we develop to control a smartphone fleet from a single computer.
@@
II.1 Sensor tests
The coordinate system of the smartphone sensors is shown in Fig. 1(a), according to the direction of the acceleration of gravity measured along the three main axes (Fig. 1(a)). Note that the $(x,y,z)$ coordinate frame of the phone forms an indirect base. The smartphones are equipped with built-in IMU sensors model ICM-42670-P fabricated by TDK InvenSense invenSense (2021).
@@ II.1 Sensor tests
II.1.1 Noise amplitude and spectrum
We first perform tests on the noise level of the smartphone sensors. We perform five-minute recordings in a quiescent environment with the accelerometer, the gyroscope and the magnetometer. The typical sampling frequency of the accelerometer is 400 Hz, with variations from a smartphone to another of about 1%. Note that for each smartphone, the sampling frequency is stable over time. A typical value of the sampling frequency is $fs=403.96\pm 0.06$ Hz (smartphone # 30). The typical acceleration error is $\sigma_{a}$ = 0.02 m s-2 in the direction perpendicular to gravity and $\sigma_{a}$ = 0.07 m s-2 along the direction of gravity. We notice that the gyroscope was sold by the constructor as a 400 Hz sampling frequency sensor, however in practice, it only operates at 50 Hz. We notice that the gyroscope measurements have indeed the same nominal sampling frequency of 400 Hz as the accelerometer. However, the gyroscope samples at 50 Hz, with repeated constant values added to artificially increase the frequency to 400 Hz. Typical error is $\sigma_{g}=0.004$ rad/s.
We interpolate the signal on a regular grid in time at a sampling frequency close to their maximum sensor sampling frequency (resp. 400 Hz for $a$ and 50 Hz for $g$ and $m$), and compute the temporal power spectrum $\hat{a}_{i}$, $\hat{g}_{i}$ and $\hat{m}_{i}$ of the acceleration, angular velocity, and magnetic field components (resp. $a_{i}$, $g_{i}$ and $m_{i}$). The power spectrum of one component for each sensor type (resp. $a_{z}$, $g_{y}$ and $m_{x}$) is shown in Fig. 2(a), (b) and (c) as a function of the frequency. The noise acceleration spectrum is flat in the entire range of tested frequency $f\in[0.1,200]$ Hz. We observe residual vibrations peaks at large frequencies. However, they may be specific to the contact between the phone and the ground, and the amplitude and locations of the peaks may not be reproducible. The noise spectrum of the gyroscope data is colored (see Fig. 2(b)) with a maximum of sensitivity around 1 Hz. The magnetometer exhibits a typical error $\sigma_{m}=1~{}\mu$T about one hundred time smaller than the Earth magnetic field. The associated noise spectrum is also colored, with a steeper slope than the gyroscope, and secondary lobes above 5 Hz.
@@ II.1.1 Noise amplitude and spectrum
II.1.2 Accelerometer sensor location
@@
We mount a smartphone on a table, rotating around a vertical axis ($Z$) at an angular frequency $\Omega$ (Fig. 1(b)). We conduct three series of experiments, with the smartphone placed with its axis $y$ (resp. $x$ and $x$) aligned with the vertical axis $e_{z}$. We vary the smartphone position with respect to the rotation axis by sliding it along the axis $x$, $y$ and $z$ respectively. The slide direction is either radial (sketches of Fig. 3(a) and (b)), or tangential (sketch of Fig. 3(c)). We then simultaneously record the accelerometer and the gyroscope signals for 60s. From a force balance in the rotating frame of reference, we expect the radial acceleration (along $e_{r}$) to be $a_{x}$ = $r\Omega^{2}$ (resp. $a_{y}$ and $a_{z}$). The acceleration $a_{x}$ vanishes when the sensor is located on the rotation axis. Note that origin of space is taken at the upper top right corner of the phone, as sketched on Fig. 1(a). Fig. 3(a,b,c) show respectively the radial accelerations $a_{x}$, $a_{y}$ and $a_{z}$ for the smartphone rotating respectively along $y$, $x$ and $x$ axes at the angular frequency $\Omega=7.7$ rad/s as a function of the smartphone locations. We indeed find a linear relationship with $r$, and we extract the position of the sensor along the three axes by interpolating the curves in $a=0$, represented by the black stars. We found $x_{S}=25.2\pm 0.1$ mm, $y_{S}=31.2\pm 0.1$ mm and $z_{S}=2.3\pm 0.1$ mm. The slope of the linear relation, $a_{x}$ = $r\Omega^{2}$ can be used to check the accuracy of the accelerometer, knowing the rotation rate $\Omega$. We found an excellent agreement (less than 1 % error) in all three direction of orientations.
Figure 4: Calibration of accelerometer and gyroscope. The measurement depends on the axis of rotation. When the smartphone rotate around the position of the gyroscope, it actually measures the angular velocity within 1% of error. When the smartphone rotates around an off-centered axis, the measured values depend on the distance $r$ to the axis of rotation. This discrepancy originates from hardware limitations. (a) 3-axis acceleration signals recorded for a rotating phone. (b) 3-axis gyroscope signals recorder in the same rotating configuration. (c) Total gyroscopic signal $g=\sqrt{g_{x}^{2}+g_{y}^{2}}$ in this rotating configuration. (d) Measured angular velocity along $y$ from $g_{x}$ signal, as a function of the position of the smartphone with respect to the rotation axis, compared to the expected value. The imposed angular speed $\Omega$ is color-coded. e) Measurement along $g_{y}$, corresponding to a rotation around $x$ that does not physically occur. This corresponds to a systematic error, which increases linearly with $r$, and increases with rotation speed. From these measurements, we infer that the minimal error is reached when the smartphone rotates around the position of the gyroscope, we find $r_{g}=31.3\pm 1$mm, different from the accelerometer location. f) Percentage of error $g_{y}/(\Omega r)$ as a function of the angular velocity $\Omega$, in cm-1.
II.1.3 Accelerometer &\&& gyroscope calibration
We now test the gyroscope accuracy using the same set up, with the accelerometer data as the reference. We rotate the phone along the $x$ axis at an angular frequency $\Omega=7.7$ rad/s, varying the position of the phone with respect to the rotation axis (sketch of Fig. 3(b)). Fig. 4(a) shows the average acceleration as a function of the relative sensor position $y-y_{S}$ to the rotation axis. The three colors corresponds to the components $a_{x}$, $a_{y}$ and $a_{z}$ of the acceleration. We recover that the sensor measures the acceleration of gravity along the $y$ axis for all sliding positions, the centrifugal force along $x$, and the $z$ component is identically zero. Fig. 4(b) shows the average angular velocity components $g_{x},g_{y}$ and $g_{z}$ measured simultaneously by the gyroscope, as a function of the relative smartphone position $y-y_{S}$. The main rotation is indeed found along $y$, and the value agrees quantitatively with less than 0.5% of error when the accelerometer sensor is on the rotation axis ($y=y_{s}$). However, we also measures a linear variation of $g_{x}$ as a function of $y-y_{S}$ and a small (quadratic) decrease of $g_{y}$. These measurements do not correspond to a true phone rotation as evidenced by the accelerometer signals, but rather to a limitation of the gyroscope sensor. In this specific case of a known rotation axis, the systematic error can be compensated, by noticing that $g=\sqrt{g_{x}^{2}+g_{y}^{2}}$ is conserved for all tested smartphone positions $y-y_{S}$ (Fig. 4(c)), and correspond to the imposed angular frequency $\Omega$. The same behaviour is observed for all three directions of rotation. This discrepancy originates from the limitation of the measurement technique implemented in the gyroscope sensor. In practice, the gyroscope is a vibrating structure gyroscope (VSG) which is based on the Coriolis effect applied on a vibrating mass. As a consequence, in a non Galilean frame of reference, the gyroscope measures a combination of inertial acceleration and Coriolis force, and the sensor reading deviates from the true angular frequency, both in magnitude and in direction. To further test the gyroscope sensor limitations, we perform additional measurements with increasing rotation rate, at four different phone locations $y-y_{S}$ for a rotation around the $x$ axis of the phone. Fig. 4(d) shows the main angular velocity component $g_{x}$ as a function of $y-y_{S}$, for different rotation rates (color coded). These measurements confirm that for small $\Omega$ (typically $\Omega<5$ rad/s), we recover $g_{x}\sim\Omega$. For larger rotation rate, we observe a significant discrepancy, which depends on the smartphone location. In particular for $|y-y_{S}|$ = 50 mm the error grows to 10% at $\Omega=10$ rad/s. The systematic error is best evidenced by the signals of $g_{y}$ (Fig. 4(e)), which should be identically zero. For all rotation speeds, the systematic error is linear with the distance $r$ between the sensor and the axis of rotation, and the slope increases with $\Omega$. From these measurements, we infer that the gyroscope sensor is located at the position $y_{g}=31.3\pm 1$ mm along the $y$ axis, very close to the accelerometer sensor. Eventually, we extract the slope $\alpha_{g}$ of the error $g_{y}=\alpha_{g}(r-r_{g})$, and we compute a relative error, $\Delta g=\alpha_{g}/\Omega$, which is measured in %$\%/cm$, as it is an error proportional to $r$. The error is represented in Fig. 4(f) as a function of $\Omega$. For a period of 1s corresponding to $\Omega=2\pi$ rad/s, we found 4% of error per centimeter  [^1] Among the sources of error, we identify (i) mechanical coupling between the $x$ and $y$ gyroscopes, which plays a symmetrical roles. It would explain why the orientation of the rotation vector is modified but the total angular velocity $\Omega$ is conserved at moderate rotation and moderate distance to the axis of rotation. The assumption of pure Coriolis force measured by $g_{i}$ fails for increasing velocity and acceleration of the gyroscope sensors, it adds at least two other sources of error : (ii) The sensor also measure inertial forces, such as centrifugal force, which is proportional to $r$. (iii) The assumption of a purely oscillatory motion imposed to the sensor by the ship is no longer valid, the velocity of the gyro with respect to the Galilean frame of reference is no longer negligible. For a rotation at constant speed, all these sources of error are proportional to the distance $r$ to the axis, but depend differently on the angular velocity $\Omega$. For oscillatory motions and a distance $r$ to the center of rotation varying in time, these systematic sources of errors would be difficult to disentangled..

### v5

@@ Abstract
Smartphones sensors are now commonly used by a worldwide audience thanks to their availability, high connectivity, and versatility. Here, we present a methodology to use a collection of smartphones, namely a fleet, as a distributed network of time-synchronized mechanical sensors. We first present the mechanical tests we develop to evaluate the smartphone sensor accuracy. We then describe how to use efficiently a distributed network of smartphones as autonomous sensors. We use a combination of an Android application hosted on each phone (Gobannos), and a server application (Phonefleet) on a controlling host to perform the tasks in parallel remotely. We implement in particular a time synchronization protocol based on UDP communication. We achieved an accuracy of the smartphone clock synchronisation of 60 microseconds. Using two test cases in realistic outdoor conditions, we eventually prove the reliability of a smartphone fleet to measure mechanical wave measurements in field conditions.
1 Introduction
@@
By November 2024, mobile subscriptions have reached 8.4 billion, including 7.14 billion for smartphones [ericsson2024ericsson]. Smartphones have become a key tool for educational and scientific research purposes thanks to their accessibility to the general public [gao2016smartphone, raju2024commentary] and to the versatility of their built-in micro-electro-mechanical system (MEMS) sensors. A typical MEMS is a highly integrated, silicon-based sensor system fabricated using semiconductor processes [maluf2002introduction]. Its low cost, compact size, lightweight design, and low power consumption make a MEMS ideal for integration into smartphones, drones, tablets, and wearable devices [daponte2013state], enabling applications such as step counting and orientation tracking [bao2004activity, naqvib2012step, davidson2016survey], gaming [hopfner2013measuring, kolakowska2020review], or navigation [li2015real, mostafa2019novel].
@@
In the context of physics education, following the COVID-19 pandemic, smartphones have gained popularity as flexible tools for providing innovative teaching methodologies, particularly at the university level [bobroff2020teaching, bouquet202061, o2021guide]. Numerous examples of tabletop physics experiments using smartphone as a measurement device can be found, such as speed of sound measurement using the embedded microphone [kasper2015stationary]; earth’s rotation estimate using the accelerometer [vandermarliere2021detect]; and light polarization analyses using ambient light sensor [monteiro2017polarization]. A comprehensive reviews of applications can be found in Organtini [organtini2021physics] or in Kuhn &\& Vogt  [kuhn2022smartphones].
@@
The versatility of smartphone MEMS sensors has also drawn significant attention from researchers across various scientific domains. The high connectivity and ubiquity of smartphones enable large-scale, real-time data collection through a geographically distributed network of interconnected devices. For public health research, investigations on the data collection and analysis of energy expenditure from smartphone users [pande2013energy] have shown that smartphone-embedded IMU (Inertial Measurement Unit) sensors such as accelerometers and gyroscopes, can accurately monitor the user’s physical activities, thus enhance the precision of individual health assessments. Smartphone inertial sensors have also been widely used for human activity recognition, where relevant motion features span a broad range of time scales, from rapid impacts to slower postural changes. Recent studies show that multiscale signal representations significantly improve activity classification while remaining compatible with the limited computational resources of smartphones. [quan2025msa, tang2022multiscale] In a review by Lee et al. [lee2018review] on smartphone applications in geoscience research, the authors highlight the growing use of smartphones in tasks traditionally carried out with specialized tools, such as geological mapping, seismic activity detection, and natural hazard assessment.
@@
However, field measurements often require multi-point data collection with sensors spatially distributed over a wide area while remaining synchronized in time. Examples of such applications include predicting solar magnetic storms and assessing their impact on the Earth’s magnetosphere [odenwald2022can], as well as performing modal analysis of civil engineering structures [gueguen2020slow], where a large number of data points are crucial for reconstructing higher-order spatial modes [cunha2006experimental, rainieri2014operational]. Distributed sensor measurements in seismology using smartphones have been introduced via real-time data collections for earthquake detection using the application MyShake [kong2016myshake]. It allows accurate localization of the earthquake epicenters and detects seismic waves propagation in real time, contributing to early warning systems [reilly2013mobile, kong2016myshake]. These applications do not rely on sophisticated, high-cost sensors but rather utilize frugal, cost-effective sensor networks, such as smartphone-based MEMS sensors fleets on a large scale, often in regions geographically inaccessible through traditional methods.
@@
Several mobile applications (PhyPhox, FizziQ[staacks2018advanced, bilgin2022stem] have been developed for individual users to facilitate sensor communication and to display data in real-time. However, there is still a notable lack of a technical way to control multiple smartphones by a single user. In contrast to single smartphone measurement that can be performed with manual interactions, distributed measurements require an automated process involving remote communication for data sampling and collection, time synchronization between smartphones, and parallel task planification. Another key challenge lies in the optimization of human-machine interaction (HMI) for a single user to control a smartphone fleet.
[…]
2 Smartphones as a multi-sensor fleet
@@
To create a fleet of autonomous, multi-physics sensors, we selected the Redmi 10A smartphone, balancing the cost and the quality of the embedded IMU sensors (accelerometer, gyroscope and magnetometer). We have also tested and benchmarked the sensors of the Fairphone 4. Still, the methodology described in this manuscript may apply to all types of smartphones. The fleet is composed of 66 smartphones, numbered from 0 to 65. We use concomitantly the accelerometer, gyroscope, magnetometer and GPS sensors. To access and utilize the smartphone sensors, we initially employed the Phyphox application [staacks2018advanced], developed at the University of Aachen. Phyphox offers an intuitive interface for acquiring data from individual sensors or from multiple sensors simultaneously, along with a standard URL communication protocol to retrieve data from the smartphone. We configured a module in Phyphox to simultaneously collect accelerometer, gyroscope, magnetometer, and GPS data at their respective maximum sampling frequencies. In a second step, we develop an Android application, named Gobannos [Gobannos_2025], better adapted to parallelized, continuous and time synchronized acquisitions without physical intervention on the phones. This section is organized as follow. Section 2.1 presents the mechanical tests performed on the smartphone sensors. Section 2.2 details the architecture we develop to control a smartphone fleet from a single computer. Some of the additional tests we performed on the smartphone fleet are presented in appendices.
@@
2.1 Sensor tests
The coordinate system of the smartphone sensors is shown in Fig. 1(a), according to the direction of the acceleration of gravity measured along the three main axes (Fig. 1(a)). Note that the $(x,y,z)$ coordinate frame of the phone forms an indirect basis. The smartphones are equipped with built-in IMU sensors model ICM-42670-P fabricated by TDK InvenSense [invenSense2021].
@@ 2.1 Sensor tests
2.1.1 Noise amplitude and spectrum
@@
We first perform tests on the noise level of the smartphone sensors. We perform five-minute recordings in a quiescent environment with the accelerometer, the gyroscope and the magnetometer. The typical sampling frequency of the accelerometer is 400 Hz, with variations from a smartphone to another of about 1%. Note that for each smartphone, the sampling frequency is stable over time. A typical value of the sampling frequency is $fs=403.96\pm 0.06$ Hz (smartphone # 30). The typical acceleration error is $\sigma_{a}$ = 0.02 m s-2 in the direction perpendicular to gravity and $\sigma_{a}$ = 0.07 m s-2 along the direction of gravity. We notice that the gyroscope was sold by the constructor as a 400 Hz sampling frequency sensor, however in practice, it only operates at 50 Hz. We notice that the gyroscope measurements have indeed the same nominal sampling frequency of 400 Hz as the accelerometer. However, the gyroscope samples at 50 Hz, with repeated constant values added to artificially increase the frequency to 400 Hz. Typical error is $\sigma_{g}=0.004$ rad/s. We tested another smartphone model, the Fairphone 4, whose accelerometer also operate around 400Hz. The effective gyroscope sampling frequency in this case was equal to the expected value of 400Hz.
[…]
@@
2.1.2 Accelerometer sensor location
@@
We mount a smartphone on a table, rotating around a vertical axis ($Z$) at an angular frequency $\Omega$ (Fig. 1(b)). We conduct three series of experiments, with the smartphone placed with its axis $y$ (resp. $x$ and $x$) aligned with the vertical axis $e_{z}$. We vary the smartphone position with respect to the rotation axis by sliding it along the axis $x$, $y$ and $z$ respectively. The slide direction is either radial (sketches of Fig. 3(a) and (b)), or tangential (sketch of Fig. 3(c)). We then simultaneously record the accelerometer and the gyroscope signals for 60s. From a force balance in the rotating frame of reference, we expect the radial acceleration (along $e_{r}$) to be $a_{x}$ = $r\Omega^{2}$ (resp. $a_{y}$ and $a_{z}$). The acceleration $a_{x}$ vanishes when the sensor is located on the rotation axis. Note that origin of space is taken at the upper top right corner of the phone, as sketched on Fig. 1(a). Fig. 3(a,b,c) show respectively the radial accelerations $a_{x}$, $a_{y}$ and $a_{z}$ for the smartphone rotating respectively along $y$, $x$ and $x$ axes at the angular frequency $\Omega=7.7$ rad/s as a function of the smartphone locations. We indeed find a linear relationship with $r$, and we extract the position of the sensor along the three axes by interpolating the curves in $a=0$, represented by the black stars. We found $x_{S}=25.2\pm 0.1$ $m$, $y_{S}=31.2\pm 0.1$ mm and $z_{S}=2.3\pm 0.1$ mm. The slope of the linear relation, $a_{x}$ = $r\Omega^{2}$ can be used to check the accuracy of the accelerometer, knowing the rotation rate $\Omega$. We found an excellent agreement (less than 1 % error) in all three direction of orientations.
2.1.3 Accelerometer &\& gyroscope calibration
The gyroscope accuracy has also been evaluated on the rotating table, using the accelerometer data as the reference. The results are detailed in A. For the Redmi 10A, we observe a significant error on the measured angular frequency, which depends on the distance between the sensor and the rotation axis. The gyroscope is a vibrating structure gyroscope (VSG) which is based on the Coriolis effect applied on a vibrating mass. As a consequence, in a non Galilean frame of reference, the gyroscope measures a combination of inertial acceleration and Coriolis force, and the sensor reading deviates from the true angular frequency, both in magnitude and in direction.